OpenStudy (anonymous):

In the below system, solve for y in the first equation. 2x + y = 11 5x - 2y = -1 CHOICES: 2x + 11 -2x + 11 2x - 11 -2x - 11 PLEASE HELP!

OpenStudy (anonymous):

\[2x - 2x + y = 11 - 2x\]

OpenStudy (anonymous):

For 2x + y = 11?

OpenStudy (anonymous):

I am lost lol

OpenStudy (anonymous):

That was a hint. You want to get y by itself. You do that by doing the same thing to both sides of the =.

OpenStudy (anonymous):

So then 5x -5x - 2y = -1-5x?

OpenStudy (anonymous):

So we can get y by itself on the left side by subtracting 2x. But that means we have to subtract 2x from the right side also. Then you can just simplify it.

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

Ok, then after that?

OpenStudy (anonymous):

If you solve this, There will be some hope. I need an example to study!

OpenStudy (anonymous):

Well, for the first one, we're basically done. Just cancel the x's.

OpenStudy (anonymous):

The second one ends up \(\frac{-2y}{-2} = \frac{-1 - 5x}{-2}\) You'll have to simplify that. Be careful with the negative signs.

OpenStudy (anonymous):

so, the answer choice would be 2x-11?

OpenStudy (anonymous):

No, that isn't what it says. 11 - 2x.

OpenStudy (anonymous):

oh, -2x + 11?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Thanks!!!!!